OpenStudy (anonymous):

Simplify the expression. Write the answer using scientific notation. 8(8.8 x 1012)

OpenStudy (unklerhaukus):

\[8\times (8.8 \times 10^{12})=70.4 \times 10^{12}=7.04 \times 10^{13}\]